# A New Era in Computational Pathology: A Survey on Foundation and Vision-Language Models

### v1

@@ I Introduction
In recent years there has been a surge of artificial intelligence (AI) based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] in CPath owing to wide adoption of digital slide scanners. As a result, large-scale curation and annotation [19, 20] of whole slide images (WSIs) has been made possible which ensured adequate data to train these AI-based models. The goal of these AI-based models is to automate and expedite the diagnosis and prognosis process of CPath. The traditional diagnosis process in CPath is time-consuming and requires experts with extensive domain knowledge. In addition, the wide variation of pathology and heterogeneity between tasks makes it difficult to come up with a unified general approach.
@@
Several research studies have addressed this lack of a unified approach and among the proposed methods, the FMs have gained a lot of attention in recent years [21, 22, 23, 24, 25, 26]. FMs leverage self-supervised learning (SSL) [27] schemes to learn a rich representation in a task-agnostic manner. Owing to self-supervised pre-training (SSPT), FMs do not require large-scale annotated data which is hard to come by in CPath. Furthermore, these models can be trained with a diverse selection of datasets containing tissue samples from different organs and associated with different cancer types, scanner types, etc. As a result, the resultant pre-trained model can easily be utilized in a wide range of downstream tasks while maintaining robustness to extreme variation in tissue samples.
@@
Figure 1: Number of publications in FMs and VLMs in pathology (from Google Scholar). The search was done with the keywords “vision-language” + “pathology” for VLMs statistics and “foundation models”+ “pathology” for FMs statistics.
@@
Figure 2: Outline of the major challenges in CPath (challenges in data collection, challenges in data annotation, lack of diverse data, large number of tasks and challenges in deep learning architectures). Several causes and consequences for each challenge are outlined in addition to how FMs and VLMs address these challenges.
@@
The impact of FMs in CPath can be amplified by integrating the power of VLMs [28] which surged in popularity after the introduction of contrastive language-image pre-training (CLIP) [29] model by OpenAI. Pathology reports, books, education videos, etc are a rich source of semantic information that can be utilized by VLMs to significantly boost performance which is not possible with vision-only models. When used in conjunction with FMs they can perform like AI pathologists capable of performing a vast array of tasks as evidenced by recent research works [25]. Even though CPath is a specialized field, there is a large increase in the number of publications focusing on FMs and VLMs as shown in Fig. 1 which indicates the future direction of this field.
@@
To appreciate the full impact of FMs and VLMs in CPath, the major challenges in CPath are outlined below. In Fig. 2 key points of these challenges are summarized. All these challenges are addressed in some form by FMs and VLMs as mentioned in Fig. 2.
@@ I-A Scope of the Review
TABLE I: Number of Surveyed articles published in top journals and conferences within 2023-2024
@@
ECCV 2
MICCAI/MICCAI Workshop 4
AAAI 1
Elsevier/Springer 3
@@
Figure 3: Visualization of the timeline of recently published work in CPath utilizing FMs and VLMs as well as multi-modal datasets. To maintain transparency we clearly annotate research articles that have been peer-reviewed and articles that are available as pre-print. Furthermore, high-impact pioneering research works published in prominent journals are highlighted. For pre-prints if there are multiple version, the latest version and the corresponding date is used.
@@ I-B Contribution and Organization
In the past few years, quite a few review articles [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have been published focusing on computational pathology. Most of these reviews include articles in the digital and computational pathology domain and the application of deep learning as a whole rather than focusing on a specific subtopic. There are a handful of papers that focus on specific architecture like MIL [32], graph-based models [33, 17], transformer-based models [13, 10], LLMs [18], etc. The contribution of this review article is mentioned below:
@@
    To the best of our knowledge, this is the first review article to summarize recent advances (most articles from 2023-2024) in foundation and vision-language models (Section III and IV). The closest peer-reviewed survey article [8] mostly provides a high-level overview without going into the details of foundation models.
@@
    A categorized and annotated timeline of the survey articles (Fig. 3) which provides a clear idea of the evolution of the FMs and VLMs in CPath.
The rest of the article is organized as follows. In section II existing multi-modal datasets are listed along with details about their source, pre-processing techniques, etc. In section III existing FMs in CPath are outlined and descriptions for the vision and vision-language pre-training schemes for these FMs are provided. In section IV an extensive list of VLMs in CPath is provided along with details about their architecture, utilized datasets and contribution. Finally, in section V the paper is concluded.
@@ II Multi-Modal Datasets in Pathology
In this section, a comprehensive summary of the existing multi-modal datasets available in CPath is provided. The modalities taken into account are vision and language. The one article that has additional modalities is CR-PathNarratives which is included because it was referenced in many other VLMs articles. All key information regarding each dataset is summarized in Table II.
@@
PathGen-1.6M [37] Image-caption pairs 1.6 million WSI and pathology report from TCGA Multi-agent collaboration and caption generation with LMMs Dataset: PathCap [38], Quilt-1M [39], OpenPath [40] Model/Framework: LLaVA-v1.5 [41], Vicuna [42] OpenCLIP [43], GPT-4 [36] PathGen-CLIP[37], PathGen-LLVA [37] 200K instruction-tuning data (Extension)
@@
CR-PathNarratives [54] WSIs with multi-modal annotations 174 annotated colorectal WSIs Proprietary data from a hospital Annotation by eight pathologists which contains ROI, voice information and behavioral trajectory information of the annotators Model/Framework: PathNarrative annotation tool [54]
@@
Dataset from MI-Zero  [62] Image-caption pairs 33,480 image-caption pairs Publicly available educational resources combined with ARCH Hand-crafted algorithms to filter out non-pathology images Dataset: ARCH MI-Zero [62]
Availability: GitHub repositories (except for ARCH [55] and PathVQA [58]; direct dataset source is linked for these two papers) with the associated paper are linked if they are open-sourced and accessible. Some repositories only provide data generation and processing code (if the data is proprietary or requires an API) and some provide direct data sources through platforms such as Hugging Face Hub, Zenodo. Other Utilized Dataset/Models: Dataset/Models/Framework used only in the data generation/ processing part is mentioned. Downstream task dataset is not mentioned for this reason.
@@ II-A Type of Datasets:
The first component is the type of data which can be broadly classified into $55$ categories. CR-PathNarratives [54] and Prov-Path [26] fall into the last category which is dissimilar to the rest of the datasets. In CR-PathNarratives, WSIs are associated with special annotated additional modalities. The additional modalities are voice and behavioral trajectory information of the pathologists who annotated the dataset. These additional modalities were collected by using the PathNarrative multimodal interactive annotation tool proposed in the same research work. The dataset Prov-Path utilizes WSI and the corresponding reports along with histopathology findings, cancer staging, genomic mutation profiles, etc collected by Providence Health System (PHS).
Figure 5: Comparison between the size of different multi-modal datasets. The size of the bubbles indicates the size of the data set and the number of elements in a data set is mentioned on the right of the bubbles. Color is used to indicate the different types.
@@
The rest of the datasets can be categorized into the remaining four categories. The image-caption/ image-text pair category (PathGen-1.5M [37], Quilt-1M [39], OpenPath [40], PathCap [38], ARCH [55]) involves a low-to-medium quality image and an associated piece of text for that image. This text can be a short caption with a description of the image or a more elaborate description. ARCH is the earliest dataset in this category that utilized Pubmed and pathology textbooks to extract the texts. PathGen-1.5 M is the latest and largest dataset in this category, but unlike other datasets in this category, the images are patches extracted from WSIs. This is in contrast to other datasets in this category like OpenPath which is constructed with Twitter posts and replies and Quilt-1M which is constructed with frames extracted from educational pathology videos. However, the curation and annotation process of this category is much easier as it can automated with hand-crafted algorithms and heuristics. But it comes with the trade-off of noisy data as due to the automated process a lot of artifacts can be present in the data.
@@
The WSI VQA/text category (PathText [34],WSI-VQA [35], PathQABench [25]) contains question-and-answer pairs or texts associated with WSIs. Comparatively, the generation process of this category is more difficult and often involves prompting LLMs to extract information or format information according to a certain template. The most common data source is the cancer genome atlas (TCGA) [19] which contains a large repository of WSI and patient report pairs. The VQA part can be of two types, close-ended question-answer pair and open-ended question-answer pair. Close-ended question-answer pairs are of a multiple-choice type or short-answer type. On the other hand, open-ended question-answer pairs contain answers that are in natural language form. Among the datasets in this category, PathQABench is unique as it contains ROI annotation of WSIs performed by expert pathologists. It has two subsets PathQABench-Public and PathQABench-Private. The former is publicly available as it was constructed with TCGA WSI and reports, and the latter was constructed with in-house data.
The next category, which is VQA (PathMMU [44], PathVQA [58], Quilt-VQA [46]) is similar to the previous category as it also contains close-ended and open-ended question-answer pairs, but the associated images are not WSIs but rather low-to-medium-quality images. Among these datasets, PathVQA is the first research to curate a pathology-specific VQA dataset. PathMMU is the latest and largest dataset in this category and it also provides explainability annotations with each answer. PathMMU utilized two previous datasets from the image-caption/image-text category OpenPath and Quilt-1M to generate a subset of the question-answer pair with images. Another category, which is the instruction-tuning dataset (Quilt-Instruct [46], PathInstruct [37], PathChatInstruct [25], extension of PathGen-1.6M [37] ) is unique kind of dataset, as this type of dataset is used to provide conversational ability to an existing multimodal model. The common workflow is that the instruction-tuning data set is applied in the last phase to fine-tune an already trained VLM. All of these datasets were created following the strategy mentioned in LLaVA [41] or LLaVa-1.5 [63]. In Fig. 5 a visual comparison of the size of each dataset in each category is provided.
II-B Source of Data:
The second component to consider is the source of data which largely dictates the third component, the annotation and pre-processing. PubMed is a common data source containing pathology images and captions/text. However, the quality of the data is not as high as that of the TCGA repository that contains WSIs and corresponding pathological reports. Other high-quality data sources that contain WSIs and pathology reports are in-house proprietary datasets. A unique data source, OpenPath, contains pathology images and texts from Twitter posts and replies associated with a large pathology community.
@@ II-B Source of Data:
This data set was supplemented by pathology-specific data from the large-scale artificial intelligence open network (LAION) data repository. Pathology textbooks and atlas are also large knowledge sources that can be used to extract image caption/text pairs. In a couple of recent research [39, 46], educational histopathology videos on YouTube are being used as the source of pathology image and text pair. However, curation of this kind of dataset requires a series of hand-crafted algorithms and many external tools.
Based on the above discussion, it is apparent that there is a trade-off between the quality and volume of the data. Sources like PathQABench are superior in terms of quality, as they were curated by an expert pathologist. However as it requires explicit manual annotation and supervision, it is much harder to scale and hence small in terms of size.
Another key point is that most of these datasets contain other datasets as one of the subsets. An example of that is Quilt-1M which contains OpenPath, PubMed and LAION as subsets in addition to the proposed Quilt dataset which is the original proposed dataset. A visualization of Quilt-1M and its subsets is provided in Fig. 6. Another such example is the PathMMU dataset (shown in Fig. 7) which contains $55$ different subsets, PubMed, SocialPath, EduContent, Atlas and PathCLS each containing data from different sources.
@@ II-C Annotation and Pre-Processing:
The third component is the annotation and pre-processing steps of data curation and generation. Among the surveyed articles, all employ a series of steps depending on the type of dataset and if the entire annotation and pre-processing pipeline is compared every article is unique. However, some specific steps are similar if the data source is the same. For example, all articles that utilize PubMed as a data source use some kind of parsing process [56] to parse and extract figures and texts. In addition, they employ light-weight classifiers and object detection architectures (YOLOv5 [60], YOLOv6 [45], YOLOv7 [51]) to distinguish between pathology and non-pathology images, detect and separate subfigures, etc. Another common approach is to prompt LLMs to format and refine captions/text or structure extracted information according to a pre-defined template. These LLMs include generalized LLMs like GPT-4, GPT-3.5, ChatGPT and also specialized LLMs like BioGPT. Another widely used strategy is using a trained CLIP-based model and using cosine similarity as a metric to classify pathology and non-pathology images.
@@ III Foundation Model
In this section, an overview of existing FMs in CPath is provided. First, the characteristics of FMs are provided to remove any ambiguity for the later sections. Next, pre-training workflow and typical pre-training schemes are mentioned.
@@ III-A Characteristics of FMs
    The first characteristic that is common to all FMs is the SSPT. The data used in the pre-training phase do not have any explicit label or annotation. Specifics of the SSL strategies are mentioned in section III-B.
@@
    In CPath, FMs are trained using large and diverse datasets that encompass tissue samples from different organs and anatomic sites. In addition, some research [64, 65] put effort into capturing diversity in terms of scanner types, magnification levels, stain types, preservation methods, etc. The idea is to capture the representation of a wide range of tissue and disease types and also making sure in downstream tasks are robust. In Fig. 8 a high-level visualization is provided highlighting different aspects of FMs.
@@
As shown in Fig. 3, the FMs are sectioned into three separate categories. One category encompasses VFMs [22] in CPath, the second category encompasses VLFMs in CPath and the last category utilizes these FMs by providing a benchmark [66, 67, 68, 69, 70], framework [71] or adapting FMs [72, 73, 74, 75, 76, 77]. Before going into the details of each category in the following section, an overview of the pre-training strategies of FMs is provided.

### v2

@@ I Introduction
In recent years there has been a surge of artificial intelligence (AI) based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] in CPath owing to wide adoption of digital slide scanners. As a result, large-scale curation and annotation [20, 21] of whole slide images (WSIs) has been made possible which ensured adequate data to train these AI-based models. The goal of these AI-based models is to automate and expedite the diagnosis and prognosis process of CPath. The traditional diagnosis process in CPath is time-consuming and requires experts with extensive domain knowledge. In addition, the wide variation of pathology and heterogeneity between tasks makes it difficult to come up with a unified general approach.
@@
Several research studies have addressed this lack of a unified approach and among the proposed methods, the FMs have gained a lot of attention in recent years[22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. FMs leverage self-supervised learning (SSL) [43] schemes to learn a rich representation in a task-agnostic manner. Owing to self-supervised pre-training (SSPT), FMs do not require large-scale annotated data which is hard to come by in CPath. Furthermore, these models can be trained with a diverse selection of datasets containing tissue samples from different organs and associated with different cancer types, scanner types, etc. Certain research works put a great emphasis on making sure this diversity is present in the pre-training dataset [26, 29]. As a result, the resultant pre-trained model can easily be utilized in a wide range of downstream tasks while maintaining robustness to extreme variation in tissue samples.
@@
Figure 1: Number of publications for FMs and VLMs in pathology (from Google Scholar). The search was done with the keywords “vision-language” + “pathology” for VLMs statistics and “foundation models”+ “pathology” for FMs statistics.
@@
Figure 2: Outline of the major challenges in CPath. Several causes and consequences for each challenge are outlined in addition to how FMs and VLMs address these challenges.
@@
The impact of FMs in CPath can be amplified by integrating the power of VLMs [44] which surged in popularity after the introduction of contrastive language-image pre-training (CLIP) [45] model by OpenAI. Pathology reports, books, educational videos, etc are a rich source of semantic information that can be utilized by VLMs to significantly boost performance which is not possible with vision-only models. When used in conjunction with FMs they can perform like AI pathologists capable of performing a vast array of tasks as evidenced by recent research works [37]. Even though CPath is a specialized field, there is a large increase in the number of publications focusing on FMs and VLMs as shown in Fig. 1 which signals the future direction of this field.
@@
To appreciate the full impact of FMs and VLMs in CPath, the major challenges in CPath are outlined in Fig. 2. In addition, the way FMs and VLMs address these challenges is also mentioned.
@@ I-A Scope of the Review
TABLE I: Number of Surveyed articles published in top journals and conferences within 2021-2024
@@
ECCV/WACV/ICCV 5
MICCAI 5
AAAI 2
Elsevier/Springer 6
@@
Figure 3: Visualization of the timeline of recently published work in CPath utilizing FMs and VLMs as well as multi-modal datasets. To maintain transparency we clearly annotate research articles that have been peer-reviewed and articles that are available as pre-print. Furthermore, high-impact pioneering research works published in prominent journals are highlighted. For pre-prints if there are multiple versions, the latest version and the corresponding date are used.
@@ I-B Contribution and Organization
In the past few years, quite a few review articles [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] have been published focusing on CPath. Most of these articles provide a review of the application of deep learning as a whole rather than focusing on a specific subtopic. There are a handful of papers that focus on specific architecture like MIL [18], graph-based models [48, 17], transformer-based models [13, 10], LLMs [19], etc. The contribution of this review article is mentioned below:
@@
    To the best of our knowledge, this is the first review article to summarize recent advances (most articles from 2023-2024) in FMs and VLMs (Section III and IV). The closest peer-reviewed survey article [8] mostly provides a high-level overview without going into the details of FMs.
@@
    A annotated timeline of the surveyed articles (Fig. 3) which provides a clear idea of the evolution of the FMs and VLMs in CPath. In addition, a project page is available (https://cpath-fms-vlms.github.io/survey/) with additional information about the surveyed articles.
The rest of the article is organized as follows: in section II existing multi-modal datasets are listed along with details about their source, pre-processing techniques, etc. In section III existing FMs in CPath are outlined and descriptions for the vision and vision-language pre-training schemes for these FMs are provided. In section IV an extensive list of VLMs in CPath is provided along with details about their architecture, utilized datasets and contribution. Finally, in section V the paper is concluded.
@@ II Multi-Modal Datasets in Pathology
In this section, a comprehensive summary of the existing multi-modal datasets available in CPath is provided. The modalities taken into account are vision and language. All key information regarding each dataset is summarized in Table II.
@@
PathGen-1.6M [51] Image-caption pairs 1.6 million WSI and pathology report from TCGA Multi-agent collaboration and caption generation with large multimodal models (LMMs) Dataset: PathCap [40], Quilt-1M [52], OpenPath [42] Model/Framework: LLaVA-v1.5 [53], Vicuna [54] OpenCLIP [55], GPT-4 [50] PathGen-CLIP[51], PathGen-LLaVA [51] 200K instruction-tuning data (Extension)
@@
Dataset from MI-Zero  [69] Image-caption pairs 33,480 image-caption pairs Publicly available educational resources combined with ARCH Hand-crafted algorithms to filter out non-pathology images Dataset: ARCH [70] MI-Zero [69]
@@
Availability: GitHub repositories (except for ARCH [70] and PathVQA [73]; direct dataset source is linked for these two papers) with the associated paper are linked if they are open-sourced and accessible. Some repositories only provide data generation and processing code (if the data is proprietary or requires an API) and some provide direct data sources through platforms such as Hugging Face Hub, Zenodo. Other Utilized Dataset/Models: Dataset/Models/Framework used only in the data generation/processing part is mentioned. The downstream task datasets are not mentioned for this reason.
@@ II-A Type of Datasets:
The first component is the type of data which can be broadly classified into $55$ categories. Prov-Path [38] fall into the last category which is dissimilar to the rest of the datasets. The dataset for Prov-Path utilizes WSI and the corresponding reports along with histopathology findings, cancer staging, genomic mutation profiles, etc collected by Providence Health System (PHS).
@@
The rest of the datasets can be sectioned into the remaining four categories. The image-caption/image-text pair category (dataset from CONCH [41], PathGen-1.5M [51], Quilt-1M [52], PathCap [40], OpenPath [42], dataset from MI-Zero [69], ARCH [70]) involves a low or medium-quality image and an associated piece of text for that image. This text can be a short caption with a description of the image or a more elaborate description. ARCH is the earliest dataset in this category that utilized PubMed and pathology textbooks to extract the texts. PathGen-1.5M is the latest dataset in this category, but unlike other datasets in this category, the images are patches extracted from WSIs.
Figure 5: Comparison between the size of different multi-modal datasets. The size of the bubbles indicates the size of the data set (For visual clarity, the scale used for bubble size is the same within a specific group, but differs between groups).
This is in contrast to other datasets in this category like OpenPath [42] which is constructed with Twitter posts and replies and Quilt-1M [52] which is constructed with frames extracted from educational pathology videos. However, the curation and annotation process of this category is much easier as it can automated with hand-crafted algorithms and heuristics. But it comes with the trade-off of noisy data as due to the automated process a lot of artifacts can be present in the data.
@@
The WSI VQA/text category (WSI-VQA [49], PathQABench [37], PathText [57]) contains question and answer pairs or texts associated with WSIs. The most common data source is the cancer genome atlas (TCGA) [20] which contains a large repository of WSI and patient report pairs. The VQA part can be of two types, close-ended question-answer pair and open-ended question-answer pair. Close-ended question-answer pairs are of a multiple-choice type or short-answer type which have pre-defined answers. On the other hand, open-ended question-answer pairs contain answers that are in natural language form. Among the datasets in this category, PathQABench is unique as it contains ROI annotation of WSIs performed by expert pathologists. It has two subsets PathQABench-Public and PathQABench-Private. The former is publicly available as it was constructed with TCGA WSI and reports, and the latter was constructed with in-house data.
@@
The next category, which is VQA (PathMMU [61], Quilt-VQA [56], PathVQA [73]) is similar to the previous category as it also contains close-ended and open-ended question-answer pairs, but the associated images are not WSIs but rather low and medium-quality images. Among these datasets, PathVQA is the first research to curate a pathology-specific VQA dataset. PathMMU is the latest and largest dataset in this category and it also provides explainability annotations with each answer. Another category, which is the instruction-tuning dataset (Quilt-Instruct [56], PathInstruct [51], PathChatInstruct [37], extension of PathGen-1.6M [51]) is a unique kind of dataset, as this type of dataset is used to provide conversational ability to an existing multimodal model. The common workflow is that the instruction-tuning data set is applied in the last phase to fine-tune an already trained VLFM or VLM. All of these datasets were created following the strategy mentioned in LLaVA [53] or LLaVA-1.5 [75]. More details about the instruction-tuning phase are provided in section III-C and section IV-C.
The comparison between dataset size is shown in Fig. 5.
II-B Source of Data:
The second component to consider is the source of data which largely dictates the third component, the annotation and pre-processing. PubMed is a common data source containing pathology images and captions/text. However, the quality of the data is not as high as that of the TCGA repository that contains WSIs and corresponding pathological reports. Other high-quality data sources that contain WSIs and pathology reports are in-house proprietary datasets (PathQABench-Private, Prov-Path). A unique data source, OpenPath, contains pathology images and texts from Twitter posts and replies associated with a large pathology community.
This data set was supplemented by pathology-specific data from the large-scale artificial intelligence open network (LAION) data repository [65]. Pathology textbooks and atlas are also large knowledge sources that can be used to extract image caption/text pairs. In a couple of recent research [52, 56], educational histopathology videos on YouTube are being used as the source of pathology image and text pair. However, curation of this kind of dataset requires a series of hand-crafted algorithms and many external tools.
Based on the above discussion, it is apparent that there is a trade-off between the quality and volume of the data.
Another key point is that most of these datasets contain other datasets as one of the subsets. An example of that is Quilt-1M which contains OpenPath, PubMed and LAION as subsets in addition to the proposed Quilt dataset. A visualization of Quilt-1M and its subsets is provided in Fig. 6. Another such example is the PathMMU dataset (shown in Fig. 7) which contains $55$ different subsets, PubMed, SocialPath, EduContent, Atlas and PathCLS each containing data from different sources.
@@ II-C Annotation and Pre-Processing:
The third component is the annotation and pre-processing steps of data curation and generation. Among the surveyed articles, all employ a series of steps depending on the type of dataset and the pre-processing pipeline is unique for every article. However, some specific steps are similar if the data source is the same. For example, all articles that utilize PubMed as a data source use some kind of parsing process [71] to parse and extract figures and texts. In addition, they employ light-weight classifiers and object detection architectures (YOLOv5 [63], YOLOv6 [62], YOLOv7 [58]) to distinguish between pathology and non-pathology images, detect and separate subfigures, etc. Another common approach is to prompt LLMs to format and refine captions/text or structure extracted information according to a pre-defined template. These LLMs include generalized LLMs like GPT-4, GPT-3.5, ChatGPT and also specialized LLMs like BioGPT. Another widely used strategy is using a trained CLIP-based model and using cosine similarity as a metric to classify pathology and non-pathology images.
@@ III Foundation Model
In this section, an overview of existing FMs in CPath is provided. First, the characteristics of FMs (section III-A) are provided to remove any ambiguity for the later sections. Next, pre-training workflow and typical pre-training schemes (section III-B) are mentioned along with the instruction-tuning phase (section III-C) which is becoming a widely adopted technique to provide FMs with conversational ability. Next, the downstream tasks and the associated downstream datasets (section III-D) for the FMs are outlined.
@@ III-A Characteristics of FMs
    The first characteristic that is common to all FMs is the SSPT. The data used in the pre-training phase do not have any explicit label or annotation.
@@
    In CPath, FMs are trained using large and diverse datasets that encompass tissue samples from different organs and anatomic sites. In addition, some research [26, 29] put effort into capturing diversity in terms of scanner types, magnification levels, stain types, preservation methods, etc. The idea is to capture the representation of a wide range of tissue and disease types. In Fig. 8 a high-level visualization is provided highlighting different aspects of FMs.
@@
As shown in Fig. 3, the FMs are sectioned into three separate categories. One category encompasses VFMs [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] in CPath, the second category encompasses VLFMs [36, 37, 38, 39, 40, 41, 42] and the last category utilizes these FMs by providing a benchmark, framework or adapting existing FMs [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87].